# Cross-Dataset Gaze Estimation by Evidential Inter-intra Fusion

### v1

@@ 4.3. Domain Generalization Results
Networks trained using only a single source domain utilize the common L1 loss function for gaze estimation, while all other settings, including the network backbone and training parameters, remain consistent with our proposed method.
Cross-dataset baseline (ETH-Gaze+Gaze360):
Network is trained by simply mixing the ETH-Gaze and Gaze360 datasets, and the network settings are the same as the single-dataset baseline.
Cross-domain methods (PureGaze, GCF):
The state-of-the-art
@@
(SOTA) gaze estimation domain generalization methods include PureGaze (Cheng et al., 2022) and GCF (Xu et al., 2023). They are trained on a single source domain (ETH-Gaze or Gaze360) using carefully designed algorithms for domain generalization.
@@
From Table 1, it can be observed that the cross-dataset training baseline (ETH-Gaze+Gaze360) achieves comparable performance with the carefully designed SOTA single-domain generalization methods on the unseen dataset. However, there is a slight error increase in the source domains. Fortunately, our proposed method EIF reduces gaze errors in both the source domains (ETH-Gaze by 0.12° and Gaze360 by 0.16°) and the unseen domains (MPIIFaceGaze by 0.60° and EyeDiap by 0.49°) compared to the baseline (ETH-Gaze+Gaze360), ultimately surpassing the best cross-domain generalization method GCF (Xu et al., 2023) (Gaze360) by a large improvement (e.g., 7.26% in EyeDiap).
4.4. Domain Adaptation Results
@@
Our previous experiments demonstrate the effectiveness of our method both in the source domain and in generalizing to unseen domains. To further validate the efficacy of our EIF framework, as done in previous domain generalization works (Xu et al., 2023; Bao and Lu, 2023), we conduct additional domain adaptation experiments.
@@
Similar with previous domain adaptation setting, we fine-tune the model with 100 images from the target domain. For our EIF, we use these data to constrain the evidence loss (Equation 14), both for all branches and their joint output after inter-fusion. For the baseline, we use these data to constrain the L1 loss. We trained each method for 100 batches, and to mitigate randomness resulting from the random sampling of the training set, we repeat the experiment 5 times and report the average results. As demonstrated in Table 2, our proposed cross-dataset gaze estimation training method significantly enhances domain adaptation performance. Notably, our EIF framework surpasses the cross-dataset baseline and achieves a reduction in angular error of 0.78° on MPIIFace and 0.29° on EyeDiap compared with the previous best single-dataset generalization method (i.e., an average improvement of 12.62%).
@@
Table 2. Comparison with SOTA methods in domain adaptation setting on 100 labeled samples. Results are shown in angle error (°).
@@
MPIIFaceGaze EyeDiap
GCF (ETH-Gaze) (Xu et al., 2023) 4.76 5.43 5.10
GCF (Gaze360) (Xu et al., 2023) 5.34 5.66 5.50
PCFGaze (ETH-Gaze) (Bao and Lu, 2023) 4.74 5.88 5.31
PCFGaze (Gaze360) (Bao and Lu, 2023) 5.41 6.39 5.90
Baseline (ETH-Gaze+Gaze360) 3.96 5.26 4.61
EIF (ETH-Gaze+Gaze360) 3.98 5.14 4.56
4.5. Ablation Study
Since our model consists of two stages, each serving a clear purpose: the first stage single-dataset training involves training a well-performing evidence learning regression branch for each dataset, while the second stage cross-dataset joint training entails fusing these branches based on evidence learning to achieve a model with superior generalization capability. Hence, for clarity, we conduct ablation experiments in two parts accordingly.
Single-dataset training: To train a well-performing single-dataset evidence learning regression branch, we primarily include two components: firstly, introducing evidential regression learning, and secondly, incorporating multiple local regressors corresponding to specific data subspaces and utilizing intra-evidential fusion to aggregate them.
Table 3. Ablation study of the single-dataset training on the ETH-Gaze Dataset.
@@ 4.5. Ablation Study
$\checkmark$ 5.02
$\checkmark$ 5.07
$\checkmark$ $\checkmark$ 4.88
As shown in Table 3, we conduct ablation studies to evaluate the effectiveness of each components mentioned above: 1) Evidence regression: To estimate uncertainty, evidential regression introduces more complex distributions to model the regression, resulting in a slight error increase of 0.05°. However, this provides the foundation for our inter-intra evidence fusion; 2) Intra-evidential fusion: Compared to using only evidential regression in gaze estimation, the approach with the intra-evidential fusion strategy significantly reduces the angle error by 0.19°, surpassing the baseline method without evidence regression by 0.14°.
Table 4. Ablation study for cross-dataset joint training. This study showcases the average angle errors (°) of the models trained on multiple source domains (ETH-Gaze+Gaze360) across the source domain test sets (ETH-Gaze and Gaze360) and the unseen domains (MPIIFaceGaze and EyeDiap).
@@
Source Unseen
$\checkmark$ 14.09 7.02
$\checkmark$ 8.25 6.92
$\checkmark$ $\checkmark$ 8.14 6.66
Cross-dataset joint training: To jointly train multiple single-dataset branches, we first adopt the inter-evidential fusion module to aggregate predictions from multiple branches. Additionally, we further designed a cross-dataset branch to integrate features from each branch. As shown in Table 4, we conduct ablation studies to evaluate the effectiveness of each module mentioned above. All the single-dataset branches used in this experiment are based on the intra-evidential fusion approach. The experimental result shows that inter-evidential fusion approach exhibits a significant reduction in errors compared to directly averaging prediction of each branch after stage 1 (Average Fusion), with reductions of 5.84° and 0.1° in the source and unseen domains, respectively. These results match the preliminary experiment and clearly verify our intuition that aggregating predictions from multiple source domains with diverse distribution is beneficial for generalization even when employing a simple average fusion strategy. However, this simply fusion approach may compromise performance in the source domain. In contrast, our inter-fusion method not only significantly boosts performance in the source domain but also further amplifies it in the target domain. Moreover, our cross-dataset branch further enhances the performance of the inter-evidential fusion, with reductions of 0.11° and 0.26° in the source and unseen domains, respectively. This result highly suggests that our designed cross-dataset branch aggregates features from each source domain branch and thus can better fuse a wide range of distributions for generalization to unseen domains.
4.6. Parameter Sensitivity Analysis
In our experiments, we consider two critical hyperparameters: the number of local regressors $G$ in single-dataset training and the MFF starting layer $K$ in cross-dataset joint training. We separately conduct experiments to analyze the impact of these hyperparameters.
Effects of the number of local regressors in single-dataset training: As shown in Table 5, we can observe that within a certain range, adjusting the $G$ has a relatively stable impact on the performance. Specifically, setting the number of local regressors to 8 achieves the best performance.
Table 5. Parameter analysis in single-dataset training on ETH-Gaze dataset.
Parameter Value Angle Error (°)
4 4.92
$G$ 8 4.88
16 4.93
@@ 4.6. Parameter Sensitivity Analysis
Effects of the layer $K$ in cross-dataset joint training: As shown in Table 6, we can see that within a certain range, adjusting the $K$ also has a relatively stable impact on the performance. In our experiment, setting $K$ to 3 achieves the best performance.
Table 6. Parameter analysis in cross-dataset training. Results are shown in angle error (°).
Param Value Dataset
Gaze360 ETH MPII EyeDiap Avg
2 11.39 4.96 6.20 7.22 7.44
$K$ 3 11.33 4.95 6.41 6.90 7.40
4 11.36 4.95 6.36 7.21 7.47
4.7. Evidence Analysis
@@
Evidential learning is crucial to our framework; we use it to achieve an inter-intra fusion, and the final evidential regression output can also provide uncertainty estimations for robust predictions. Therefore, in this section, we offer a visual analysis of the fusion process and uncertainty predictions to validate the effectiveness of evidential learning in gaze estimation.
Inter-intra fusion visual analysis: Based on Equation 14, among the four output values of the evidential regression, $\delta$ represents the predicted gaze. Furthermore, according to Equation 11, during the evidential fusion process, normalized $\gamma$ is employed as weights to aggregate $\delta$ accordingly. Hence, to further delve into our inter-intra evidential fusion module, we compute the mean normalized $\delta$ values spanning multiple local regressors or branches.
As depicted in Fig. 4(a), the visualization showcases the output on the yaw values of the test set corresponding to the single-dataset branch in ETH-Gaze. It is observable that the diagonal region possesses the highest values, indicating our intra-fusion module enables samples within each gaze group to find their corresponding local regressors effectively. Moreover, it is noted that the weights are not solely concentrated on a single diagonal but are diffused around the diagonal line. This distribution is attributed to our adoption of a density overlap strategy for gaze group segmentation (Equation 2), which allocates multiple experts to one sample in overlapping areas, thereby enhancing robustness. Conversely, due to the lack of overlap in the boundary gaze group areas (0 and 7), the weights are relatively more focused compared to other regions.
As shown in Fig. 4(b), the experiment is conducted on the complete EIF structure. It demonstrates that all source domain test sets have the highest weights on their respective single-dataset branches. On the other hand, in the unseen domains, the weights are distributed evenly across various branches. Notably, EyeDiap has the highest weight on the cross-dataset branch, while MPIIFaceGaze shows the highest weight on the ETH-Gaze branch. The possible reason is that the single-dataset branch learned on the ETH-Gaze dataset demonstrates a strong generalization capability on MPIIFaceGaze. Previous domain generalization methods in Table 1 have also shown that models trained on ETH-Gaze perform well on MPIIFaceGaze, which is consistent with this observation and verifies the effectiveness of our evidential learning fusion strategy.
Figure 4. Visualizing Weights During Inter-Intra Fusion.
Uncertainty Estimation Analysis: Our model can simultaneously provide epistemic uncertainty and aleatoric uncertainty. Epistemic uncertainty reflects the model’s uncertainty about the underlying structure of the data. It arises from the model’s lack of knowledge or exposure to certain regions of the input space during training. As shown in Fig. 5, the epistemic uncertainty of the unseen EyeDiap dataset is relatively large in our EIF structure’s final output, while that of the source domains and MPIIFaceGaze is smaller. The smaller uncertainty in MPIIFaceGaze may be due to its similarity to a subset of the source domain datasets.
Figure 5. The epistemic uncertainty of the dataset.
The aleatoric uncertainty of the data involves the inherent uncertainty of the data itself, which cannot be eliminated even with a sufficiently large dataset. This uncertainty arises from factors such as low image quality, including reasons like image blur and occlusion of the eyes in our task.
Figure 6. The aleatoric uncertainty of two source datasets
As depicted in Fig. 6, we showcase the aleatoric uncertainty output of our EIF in the source domain test sets, where ETH-Gaze is collected on a per-user basis, and Gaze360 is collected on a per-scene basis. Therefore, we respectively output the maximum and minimum values of the estimated aleatoric uncertainty for each subject or scene on ETH-Gaze and Gaze360. It can be observed that the aleatoric uncertainty output by our model effectively reflects the data quality in the source domain. More visualization results can be found in the appendix.
4.8. Computational Cost
Despite encompassing multiple branches, thanks to our two-stage learning strategy, our model can be trained swiftly to adapt to new scenarios. Once we have multiple trained single-dataset branches, our model requires only a minimal number of epochs in stage 2 to rapidly adjust to changes in the dataset, such as additions or deletions. Specifically, in our experiments, the training of stage 2 takes only 40 minutes on two NVIDIA TITAN XP GPU, whereas the baseline needs to start from scratch and mixing multiple datasets, which requires 5 hours for training.
5. conclusion
@@ 4.8. Computational Cost
Training with diverse source domain data can significantly improve the model’s generalization ability in unseen domains. However, simple mixing may lead to adverse effects among datasets with inherent distribution shifts, thereby increasing their domain-specific errors. Our paper addresses these challenges by proposing a two-stage training approach. In the first stage, we utilize single dataset branches tailored to each source domain. Additionally, in stage 2, we conduct joint training of multiple single-dataset branches and introduce a cross-dataset branch to further aggregate the features extracted from the source domain branches. Considering the non-stationary nature of gaze behavior, we further divide individual source domain data into multiple subsets within each branch, each designed with a local regressor. We employ an evidential fusion strategy to integrate the predictions from various branches and local regressors. The experiments demonstrate the effectiveness of our method in both the source and unseen domains.

### v2

@@ 4.3 Domain Generalization Results
Single-dataset baselines are trained on a single source domain using the standard L1 loss function for gaze estimation, while other network settings remain consistent with those of our proposed method.
Cross-dataset baselines are trained on multi-source using two different mixing strategies: Simple Mixing, where the datasets are directly concatenated, and Balanced Mixing, which, similar to our Stage 2 method, uses balanced sampling. The network and training settings are kept identical to those in the single-dataset baseline.
Multi-source domain generalization methods aim to explore combining multiple source datasets to enhance performance in the target domain, making them relevant to our scenario. We implement two representative regression methods: GMDG [32] and BNE [30], with training settings consistent with the balanced mixing cross-dataset baseline. Model details are provided in the supplementary.
@@
Cross-domain gaze estimation are trained on a single source domain using carefully designed algorithms. We report two state-of-the-art approaches [11, 4] for comparison.
@@
From Table 1, it can be observed that the cross-dataset baseline achieves comparable performance with the carefully designed SOTA single-domain generalization methods on the target dataset. However, there is a slight error increase in the source domains. Multi-source domain generalization fails to bring improvements and may even slightly degrade performance. In contrast, our proposed method EIF reduces gaze errors in both the source domains (ETH-Gaze by 0.48° and Gaze360 by 0.30°) and the target domains (by 0.22°) compared to the best cross-dataset baseline (balanced mixing), ultimately surpassing the best cross-domain generalization method AGG [4] (ETH-Gaze) by a large improvement (1.14° in source domain and 0.86° in target domains).
4.4 Domain Adaptation Results
@@
Our previous experiments demonstrate the effectiveness of our method both in the source domain and in generalizing to target domains. To further validate the efficacy of our EIF framework, as done in previous domain generalization works [11, 3], we conduct domain adaptation experiments.
@@
Similar to the previous domain adaptation setting [3], we fine-tune the model using 100 images from the target domain over 100 batches. For our EIF, we use these data to constrain the evidence loss (Eq. 15), both for all branches and their joint output after inter-fusion. To mitigate randomness from selecting the training set, we repeat the experiment 5 times and report the average results.
As shown in Table 2, our EIF framework significantly improves performance, reducing the error by 0.83° compared to the second best method, PCFGaze (ETH-Gaze).
@@
Table 2: Comparison with SOTA methods in domain adaptation on 100 labeled samples. Results are shown in angle error (°).
@@
MPII EyeDiap
PureGaze (ETH-Gaze) [11] 5.30 6.42 5.86
PureGaze (Gaze360) [11] 5.20 7.36 6.28
PCFGaze (ETH-Gaze) [3] 4.74 5.88 5.31
PCFGaze (Gaze360) [3] 5.41 6.39 5.90
EIF (ETH-Gaze+Gaze360) 3.91 5.04 4.48
4.5 Ablation Study
Our model consists of two stages: single-dataset training to train a well-performing evidence learning regression branch for each dataset, and cross-dataset joint training to fuse these branches for improving cross-dataset generalization. To verify, we conduct ablation experiments in two parts.
Single-dataset training. It primarily includes two components: evidential regression learning and the use of multiple local regressors with intra-evidential fusion to aggregate them.
Table 3: Ablation study of the single-dataset training on the Gaze360 Dataset.
@@ 4.5 Ablation Study
$\checkmark$ 4.55
$\checkmark$ 4.59
$\checkmark$ $\checkmark$ 4.35
As shown in Table 3, we evaluate the components through ablation studies: 1) Evidence regression introduces complex distributions to estimate uncertainty, slightly increasing the error by 0.04°, but providing the foundation for inter-intra evidence fusion. 2) Intra-evidential fusion reduces angle error by 0.24° compared to evidential regression alone, surpassing the baseline method without evidence regression by 0.20°.
Table 4: Ablation study for cross-dataset joint training. It showcases the average angle errors (°) of the models trained on multiple source domains (ETH+G360) across the source domain test sets (ETH and G360) and the target domains (MPII and EyeDiap).
@@
Source target
$\checkmark$ 11.32 6.86
$\checkmark$ 7.47 6.67
$\checkmark$ $\checkmark$ 7.44 6.23
Cross-dataset joint training. To jointly train multiple single-dataset branches, we use the inter-evidential fusion module to aggregate predictions and design a cross-dataset branch to integrate features. As shown in Table 4, ablation studies show that: 1) Inter-evidential fusion reduces errors by 3.85° and 0.19° in the source and target domains, respectively, compared to average fusion, which performs well in the target domain but worse in the source domain; 2) Cross-dataset branch further reduces errors by 0.03° and 0.44° in the source and target domains, respectively. These results demonstrate that the cross-dataset branch effectively aggregates features from each source domain, particularly enhancing generalization to the target domain.
4.6 Parameter Sensitivity Analysis
@@
We consider two critical hyperparameters: 1) The number of local regressors $G$ in single-dataset training: As shown in Fig. 4(a), adjusting $G$ within a certain range has a stable impact on performance, with $G=8$ yielding the best results. 2) The MFF starting layer $K$ in cross-dataset joint training: As shown in Fig. 4(b), adjusting $K$ also stabilizes performance within a certain range, with $K=3$ achieving the optimal performance.
Figure 4: Parameter Sensitivity Analysis.
4.7 Evidence Analysis
@@
Evidential learning is crucial to achieve an inter-intra fusion in our framework. Besides, the final evidential regression output can also provide uncertainty estimations for robust predictions. Thus, in this section, we offer a visual analysis of the fusion process and uncertainty prediction to validate the effectiveness of evidential learning in gaze estimation.
Inter-intra fusion visual analysis.
Figure 5: Visualizing Weights During Inter-Intra Fusion.
According to Eq.12, during the evidential fusion process, normalized $\gamma$ is used as weights to aggregate the outputs $\delta$. Therefore, we present the normalized $\gamma$ values for the intra-fusion of the single-dataset branch in ETH-Gaze and the inter-fusion of EIF. In Fig. 5(a), the highest values appear in the diagonal region, indicating that the intra-fusion module effectively matches samples within each gaze group to their corresponding local regressors. In Fig. 5(b), source domain test sets assign the highest weights to their respective single-dataset branches, while target domains distribute the weights more evenly across branches, which aligns with our expectations.
Uncertainty estimation analysis.
Figure 6: Visualizing Epistemic and Aleatoric Uncertainty.
We visualize the epistemic and aleatoric uncertainty output by EIF in Fig. 6. To mitigate gaze range differences across datasets, epistemic uncertainty is computed only for gaze within ±20° for all datasets. In Fig. 6(a), epistemic uncertainty is smaller in the source domain (ETH+G360) compared to the target domain (EyeDiap+MPII), with MPII showing lower uncertainty than EyeDiap, likely due to its closer similarity to the source domain. Fig. 6(b) shows that images with higher aleatoric uncertainty indeed appear more ambiguous. More details are provided in the supplementary.
5 Conclusion and Discussion
@@
Training with diverse source domain data enhances model generalization in target domains. However, simple mixing can worsen domain-specific errors due to distribution shifts. To address this, we propose a two-stage training approach. In stage 1, we use single dataset branches for each source domain. In stage 2, we combine these branches with a cross-dataset branch to aggregate features. Given the non-stationary nature of gaze behavior, we divide each source domain into subsets with local regressors. We apply evidential fusion to integrate predictions from all branches and regressors. Experiments show the effectiveness of our approach in both source and target domains.
Limitation. While our EIF method has demonstrated significant improvements in cross-dataset training with two commonly used, highly divergent source domains, a limitation is the lack of consideration for handling an increasing number of source domain datasets with similarities. Specifically, using a single branch shared by multiple similar domains could help save resources. To address this, we plan to explore clustering similar datasets and allowing them to share a single branch, thus improving the scalability.
Supplementary Material
In this supplementary material, we provide further details on our proposed methods and experiments. The first section offers a detailed introduction to the dataset used in our cross-dataset gaze estimation setting, including a visual display of the different dataset collection environments, along with example samples and their label distributions. This section highlights the differences between the datasets and underscores the necessity of our work in cross-dataset gaze estimation. The second section provides an overview of the representative multi-source domain generalization methods we selected for comparison. The third section presents additional visualization results related to aleatoric uncertainty. Together, these sections provide a comprehensive illustration of our methodology.
1 Detailed Dataset Descriptions
In this section, we will provide a detailed introduction to the gaze estimation datasets. This explanation will help further clarify the motivation behind our proposed cross-dataset gaze estimation setting and the design of our method.
Figure 7: Visual display of the different dataset collection environments, along with data examples and their label distributions.
The experiments utilize four widely used gaze estimation datasets: EyeDiap [14], MPIIFaceGaze [40], Gaze360 [20], and ETH-XGaze [41]. For a fair comparison, the data partitioning and preprocessing techniques for these datasets are kept consistent with prior studies, as outlined in [10]. As illustrated in Fig 7, these datasets were collected under different setups and environments, resulting in substantial differences in the data samples and label distributions. Below, we provide a detailed description of each dataset.
ETH-Gaze: This dataset is captured in a laboratory environment using high-resolution cameras and consists of 756k training images from 80 subjects, exhibiting a wide range of gaze label variations. Since the ETH-Gaze dataset does not include gaze target annotations for the test set, we redefine the dataset splits following the approach in [10]. Specifically, we split the original training set into two subsets to use as training and testing sets, as done in previous studies [9].
Gaze360: This dataset is collected in uncontrolled environments using a 360-degree camera. It includes 84k training images from 54 subjects and 16k testing images from 15 subjects, featuring a wide variety of gaze label distributions.
MPIIFaceGaze: This dataset is captured during natural laptop use, utilizing the laptop camera for image capture. It consists of 45k images from 15 users, containing a limited set of gaze labels, specifically only frontal gaze data.
EyeDiap: This dataset is collected in a controlled laboratory setting, where participants are asked to gaze at two types of visual targets: screen targets and a 3D floating ball. It includes 16k images from 16 users, with a restricted set of gaze labels, primarily focused on frontal gaze data.
As observed above, inherent distribution shifts exist between datasets, including variations in image distributions (e.g., collection environments, resolution, and head pose) and gaze label ranges. These disparities make it challenging to merge multiple datasets, highlighting the need for research in cross-dataset gaze estimation. From the perspective of method design, due to distribution shifts among datasets, maintaining separate models for each dataset is often necessary in practice. This motivates our two-stage approach. In the first stage, each single-dataset branch is trained separately. Considering the broad label distribution and non-stationary gazing processes within each dataset, multiple local regressors are employed for each individual dataset, enhancing the performance of with-in domains. In the second stage, building upon the multiple single-dataset branches retained in the production environment, the model enables rapid generalization to new dataset combinations.
2 Overview of Multi-Source DG
In this section, we provide a detailed introduction of the two Multi-source Domain Generalization methods, GMDG [32] and BNE [30], which are chosen for performance comparison with our approach.
GMDG: This method is a recent and highly representative approach that introduces a novel and general learning objective, aiming to interpret and analyze the core principles underlying most existing strategies in multi-domain generalization (mDG). Unlike traditional domain generalization methods, which primarily focus on classification tasks, GMDG offers a more generalized framework capable of addressing a wide range of tasks, including regression tasks like ours.
The core of GMDG is a general learning objective that combines four optimization goals, so the loss function is designed as a weighted sum of these four terms, each targeting a specific aspect of the optimization process:
  • GAim1: This term encourages alignment across domains by minimizing the entropy of the joint distribution of features, $P(\varphi(X),\psi(Y))$. The loss for this term is:
    $L_{A1}=\sum_{i=1}^{n}\left(\log|\Sigma_{i}|+\|\bar{\mu}-\mu_{i}\|^{2}_{\Sigma_{i}^{-1}}\right),$ (21)
    where $\Sigma_{i}$ is the covariance matrix of the feature distribution in domain $i$, and $\mu_{i}$ is the mean feature vector.
  • GAim2: This term further enhances alignment by minimizing the entropy between different arrangements of features, $P(\psi(Y),\varphi(X))$ and $P(Y,\psi(Y))$. The corresponding loss is:
    $L_{A2}=H(P(\psi(Y),\varphi(X)))+H(P(Y,\psi(Y))),$ (22)
    where $H$ denotes the entropy of the distribution.
  • GReg1: This regularization term minimizes the Kullback-Leibler (KL) divergence between the joint distribution $P(\varphi(X),\psi(Y))$ and a prior distribution $O$ (ImageNet pretrained ResNet-18), which is introduced to prevent overfitting. The loss for this term is:
    $L_{R1}=D_{\text{KL}}(P(\varphi(X),\psi(Y))\|O),$ (23)
    where $D_{\text{KL}}$ is the Kullback-Leibler divergence.
  • GReg2: This term suppresses invalid causal relationships by minimizing the Conditional Feature Shift. The corresponding loss is:
    $L_{R2}=\|\Sigma_{XY}\Sigma_{YY}^{-1}\Sigma_{YX}\|^{2},$ (24)
    where $\Sigma_{XY}$, $\Sigma_{YY}$, and $\Sigma_{YX}$ are the covariances between different feature distributions.
The overall loss function is the weighted combination of these individual loss terms:
$\small L(C,\varphi,\psi)=v_{A1}L_{A1}+v_{A2}L_{A2}+v_{R1}L_{R1}+v_{R2}L_{R2},$ (25)
where we select the loss weights $v_{A1}=0.001$, $v_{A2}=1$, $v_{R1}=0.01$, and $v_{R2}=0.0001$, based on the depth estimation task, which is a similar regression task to ours, as described in the original paper [32].
BNE: This method is a representative approach for ensemble learning-based multi-source domain generalization, which are compared for two reasons: first, our multi-branch evidential fusion approach can also be seen as an ensemble learning method; second, it shares similarities with the related work on cross-dataset training in autonomous driving, which also addresses inherent distribution shifts using shared BN layers [34].
In this approach, Batch Normalization (BN) parameters $(\mu_{d},\sigma_{d}^{2})$ are maintained separately for each domain $d$, while other model parameters are shared across domains. During inference, the final prediction $p_{t}$ for sample $t$ is computed as a weighted combination of predictions $p_{d}^{t}$ from each domain-specific model. The weight $w_{d}^{t}$ is based on the similarity between the test instance statistics $r_{t}^{l}=(\mu_{t}^{l},\sigma_{t}^{l})$ at layer $l$ and the accumulated domain statistics $e_{d}^{l}=(\mu_{d}^{l},\sigma_{d}^{l})$, with similarity measured using the 2-Wasserstein distance $\mathcal{W}$:
$w_{d}^{t}=\frac{1}{\sum_{l\in\mathcal{B}}\mathcal{W}(e_{d}^{l},r_{t}^{l})},$ (26)
where $\mathcal{B}$ denotes the set of layers used for the distance computation.
Finally, the prediction $p_{t}$ is calculated as the weighted average of the domain-specific predictions:
$p_{t}=\frac{\sum_{d\in\mathcal{D}}w_{d}^{t}p_{d}^{t}}{\sum_{d\in\mathcal{D}}w_{d}^{t}},$ (27)
where $p_{d}^{t}$ is the prediction from the domain-specific model for domain $d$, and $w_{d}^{t}$ is the computed weight based on domain similarity.
Figure 8: The aleatoric uncertainty of ETH-Gaze test set.
Figure 9: The aleatoric uncertainty of Gaze360 test set.
3 Aleatoric Uncertainty Visualization
As shown in Fig. 8 and Fig. 9, we provide additional visualizations of aleatoric uncertainty for the test sets of two source domain datasets. To facilitate comparison, we visualize the top 3 highest and lowest aleatoric uncertainty values at the user level in the ETH-Gaze dataset and at the scene level in the Gaze360 dataset, since the ETH-Gaze dataset is collected at the individual level, while Gaze360 is collected at the scene level. The results demonstrate that images with occluded, blurred, or invisible eyes exhibit relatively higher uncertainty, showing that our method can accurately assess aleatoric uncertainty.